How many atoms are there in 0.0445 g of carbon
1
Expert's answer
2020-10-06T09:19:37-0400

Question # 137005

How many atoms are there in 0.0445 g of carbon?

Answer

Mol of carbon = 0.0445/12

                                     = 0.003708333 mol

             1 mol of carbon→ 6.022×1023 atoms

             0.003708333 mol →? Atoms

             =0.003708333×6.022×1023 atoms

             = 2.233158333×1021 atoms

             = 2.23×1021 atoms